# library.bib

### v1

@@ A.4 Hardware Specifications
The 2$\times$2 DDU-Net101dp and UNet101dp models, trained on the large data-driven dataset of 101 samples, were trained and evaluated on the DelftBlue cluster [DHPC2024], using NVIDIA V100 GPUs with 32 GB memory. All training was conducted using PyTorch 2.1.0 with CUDA 11.6 acceleration.
Training and evaluation of the LGCNN model were performed on a single NVIDIA A100-SXM4 GPU. Data generation was carried out on a dual-socket system equipped with AMD EPYC 9274F CPUs.
@@ A.5 Glossary
A list of the most relevant physical properties used in our paper is provided in Table 17.
@@
$s$ central streamlines after all $i$
$s_{o}$ streamlines with transversal offset to $i$

### v2

@@ A.4 Hardware Specifications
The vanilla models that were trained on the large data-driven dataset of 101 samples were trained and evaluated on a server using NVIDIA V100 GPUs with 32 GB memory. All training was conducted using PyTorch 2.1.0 with CUDA 11.6 acceleration.
Training and evaluation of all other models (dilated UNet, LGCNN single steps, full pipeline and ablation studies, FNOs, PINNs) were performed on a single NVIDIA A100-SXM4 GPU. Data generation was carried out on a dual-socket system equipped with AMD EPYC 9274F CPUs.
@@ A.5 Glossary
A list of the most relevant physical properties used in our paper is provided in Table˜13.
@@
$s$ central streamlines, starting at all $i$
$s_{o}$ streamlines starting with $\pm$ 10 cells transversal offset compared to $s$